# A Symplectic Proof of the Quantum Singleton Bound

### v1

@@ Abstract
We present a symplectic linear algebraic proof of the Quantum Singleton Bound for stabiliser quantum error–correcting codes, together with a complete Lean4 formalisation of the linear–algebraic argument. The argument is formulated in terms of a finite dimensional symplectic vector space model of Pauli operators, distance based erasure correctability, and the cleaning lemma.
1 Motivation
The Quantum Singleton Bound is a basic constraint on the parameters of a stabiliser code, but it is often proved using entropy inequalities and other information–theoretic tools. For stabiliser codes, however, the underlying structure is symplectic linear algebra: Paulis form a finite-dimensional symplectic space, commutation is bilinear, and erasure correctability is captured by the inclusion $S^{\perp}\cap V_{E}\subseteq S$. This viewpoint yields a proof that is conceptually aligned with the stabiliser formalism and isolates the algebraic ingredients (distance $\Rightarrow$ correctable erasures, plus the cleaning lemma).
A further benefit is mechanisation: the linear-algebraic argument is well-suited to formal verification, and we have formalised it in Lean4. In contrast, entropy-based proofs typically require substantially heavier analytic infrastructure, making them harder to machine-check in a lightweight way.
@@ 2 Contributions
    give a symplectic linear–algebraic proof of the Quantum Singleton Bound for $[[n,k,d]]$ stabiliser codes, deriving $k+2(d-1)\leq n$ from distance-based erasure correctability and the cleaning lemma via a dimension argument;
@@
    accompany the proof with a Lean4 formalisation of the main linear–algebraic development.
3 Introduction
The Quantum Singleton Bound is a fundamental constraint on the parameters of a quantum error– correcting code. For an $[[n,k,d]]$ stabiliser code it states
$k+2(d-1)\leq n.$
We prove this bound using the symplectic stabiliser formalism together with two key structural inputs:
  1. (i)
    distance implies correctability of erasures of size at most d1d-1;
  2. (ii)
    the cleaning lemma and its dimension identity $g(M)+g(M^{c})=2k$.
@@ 4 Conventions and symplectic model
with elements written $v=(v_{X},v_{Z})$.
@@
Fix $i$. Pairing with $(0,e_{i})$ forces $u_{X}(i)=0$, and pairing with $(e_{i},0)$ forces $u_{Z}(i)=0$. Hence all coordinates vanish. ∎
@@ 5 Support, restriction, and distance
Restriction to $C$ yields an isomorphism $V_{C}\cong\mathbb{F}^{|C|}\times\mathbb{F}^{|C|}$. ∎
@@
For $E\subseteq[n]$, define $r_{E}:V\to V_{E}$ by zeroing coordinates outside $E$.
@@
A stabiliser code $Q$ is specified by an isotropic subspace $S\subseteq V$ with $\dim(S)=n-k$. Define the (Pauli) distance
@@
Nondegeneracy gives $\dim(S)+\dim(S^{\perp})=2n$, hence $\dim(S^{\perp})=n+k$ and $\dim(S^{\perp}/S)=2k$. ∎
@@ 6 Erasure correctability
In the stabiliser setting this condition is equivalent to the standard Knill-Laflamme error-correction condition for erasures supported on $E$ [NC00].
@@
If $v\in S^{\perp}\cap V_{E}$ and $v\notin S$, then $v\in S^{\perp}\setminus S$ with $\mathrm{wt}(v)\leq|E|\leq d-1$, contradicting the definition of $d$. ∎
@@ 7 Cleaning lemma and the dimension identity
This counts the number of independent logical Pauli operators (modulo stabilisers) that can be supported on $M$.
@@
The following argument is a linear–algebraic reformulation of the dimension-counting proof appearing in Preskill’s notes [PRE22].
Let $P\cong V$ denote the abelianised Pauli group viewed as a 2n2n–dimensional $\mathbb{F}$–vector space with the same symplectic form. Fix $M\subseteq[n]$ and write $M^{c}$ for its complement. Let $P_{M}:=V_{M}$ and $P_{M^{c}}:=V_{M^{c}}$.
Define the stabiliser–restriction subspaces
@@
Indeed, if a nontrivial linear combination of elements of $S_{M}\oplus S_{0}$ restricts to zero on $M$, then it is supported on $M^{c}$, hence lies in $S\cap P_{M^{c}}=S_{M^{c}}$, which is impossible unless the combination was trivial.
Now compute $\dim(S^{\perp}\cap P_{M})$. Since $P_{M}$ is orthogonal to $S_{M^{c}}$ (disjoint support), the constraints on commuting with $S$ inside $P_{M}$ are exactly orthogonality to $r_{M}(S_{M}\oplus S_{0})=S_{M}\oplus r_{M}(S_{0})$. Therefore,
@@
It is worthwhile noting that the identity $g(M)+g(M^{c})=2k$ does not imply that each logical Pauli operator is supportable on exactly one of $M$ or $M^{c}$.
Lemma 6 (Cleaning lemma (support on the complement)).
@@
If $M$ is correctable then $S^{\perp}\cap V_{M}\subseteq S$, hence $g(M)=0$. By the dimension identity, $g(M^{c})=2k=\dim(S^{\perp}/S)$. ∎
8 A symplectic consequence of correctable erasure
@@
Set $D:=A^{c}=B\cup C$. Since $A$ is correctable we have $g(A)=0$, hence by the cleaning dimension identity
@@
so the quotient space
$L:=(S^{\perp}\cap V_{D})/(S\cap V_{D})$
has dimension $\dim(L)=2k$.
@@
We claim $\bar{r}_{C}$ is injective. Suppose $\bar{r}_{C}([v])=0$. Then $r_{C}(v)\in r_{C}(W)$, so there exists $w\in W$ with $r_{C}(v)=r_{C}(w)$. Hence $r_{C}(v-w)=0$, so vwv-w is supported entirely on $B$, i.e. $v-w\in V_{B}$. Also $v-w\in S^{\perp}$ (since $v\in S^{\perp}$ and $w\in S$). Thus $v-w\in S^{\perp}\cap V_{B}$, and because $B$ is correctable we get $v-w\in S$. Since $w\in S$ as well, it follows that $v\in S$, hence $[v]=0$ in $L$.
@@
9 Quantum Singleton bound
@@
Choose disjoint $A,B\subseteq[n]$ with |A|=|B|=d1|A|=|B|=d-1 and let $C=[n]\setminus(A\cup B)$. By distance–correctability, erasures of $A$ and $B$ are correctable. Lemma 7 yields
@@
If instead n<2(d1)n<2(d-1), then no such disjoint subsets exist. In this regime the inequality
$k\leq n-2(d-1)$
has a non-positive right-hand side, and since $k\geq 0$ for any stabiliser code, the bound is automatically satisfied. Hence the argument above covers all cases.
10 Conclusion
We derived the Quantum Singleton Bound from (i) symplectic stabiliser linear algebra, (ii) distance–based erasure correctability, and (iii) a dimension–counting proof of the cleaning lemma following Preskill. The final inequality follows from a purely symplectic dimension argument using correctability of two disjoint erasures.

### v2

@@ Abstract
We present a symplectic linear-algebraic proof of the Quantum Singleton Bound for stabiliser quantum error-correcting codes, together with a complete Lean4 formalisation of the linear-algebraic argument. The proof is formulated in the language of finite-dimensional symplectic vector spaces modelling Pauli operators and relies on distance-based erasure correctability and the cleaning lemma. Using a dimension-counting argument within the symplectic stabiliser framework, we derive the bound $k+2(d-1)\leq n$ for any $[[n,k,d]]$ stabiliser code. This approach isolates the algebraic structure underlying the bound and avoids the heavier analytic machinery that appears in entropy-based proofs, while remaining well-suited to formal verification.
1 Introduction
Quantum error-correcting codes are essential for protecting quantum information against decoherence and operational noise, a prerequisite for scalable fault-tolerant quantum computation [KL97, GOT97, CS96, STE96]. Among the most fundamental constraints on such codes is the Quantum Singleton Bound: for an $[[n,k,d]]$ stabiliser code,
$k+2(d-1)\leq n.$
This is the quantum analogue of the classical Singleton bound $k+d-1\leq n$ for linear codes over finite fields [SIN64, MS77], and it places a hard limit on the trade-off between the number of encoded logical qubits $k$, the code distance $d$, and the block length $n$. Codes that achieve equality are called quantum maximum-distance separable (MDS) codes and have been the subject of extensive study [GGB99, RAI98].
The bound was first established by Knill and Laflamme [KL97] and independently by Bennett, DiVincenzo, Smolin, and Wootters [BDS+96] in the context of general quantum error-correcting codes. Their arguments rely on properties of quantum states and channels, and the standard textbook proof [NC10] proceeds via von Neumann entropy inequalities: one shows that if a code can correct erasures of d1d-1 qubits, then the encoded information must be recoverable from the remaining n(d1)n-(d-1) qubits, and an entropy argument constrains how much information can be stored. More recently, Grassl, Huber, and Winter [GHW22] gave a streamlined entropic proof and extended the technique to entanglement-assisted and catalytic codes.
For stabiliser codes, however, the underlying mathematical structure is fundamentally algebraic, not information-theoretic. The $n$-qudit Pauli group (modulo phases) is naturally identified with a $2n$-dimensional vector space over a prime field $\mathbb{F}_{p}$, equipped with a symplectic bilinear form that encodes commutation relations [CRS+97, CRS+98, GOT97]. A stabiliser code is then specified by an isotropic subspace of this symplectic space, and concepts such as error detection, correction, and code distance admit clean algebraic formulations. A natural question is therefore whether the Quantum Singleton Bound can be derived purely within this symplectic framework, without invoking entropy or the theory of quantum channels.
In this paper we give such a proof. The argument combines three ingredients, each of which is standard in the stabiliser literature but which together yield a self-contained and concise derivation:
  1. (i)
    Distance implies erasure correctability. If a stabiliser code has distance $d$, then any set of at most d1d-1 qudit positions forms a correctable erasure, which in the symplectic model is captured by the inclusion $S^{\perp}\cap V_{E}\subseteq S$.
  2. (ii)
    The cleaning lemma. For any partition of qudit positions into a set $M$ and its complement $M^{c}$, the dimensions of the spaces of “supportable” logical operators on $M$ and $M^{c}$ satisfy $g(M)+g(M^{c})=2k$. This identity, due in its dimension-counting form to Preskill [PRE22], captures how logical information can be “cleaned” off a correctable region.
  3. (iii)
    A dimension argument. Given two disjoint correctable sets $A$ and $B$ with |A|=|B|=d1|A|=|B|=d-1, the cleaning lemma forces all $2k$ independent logical operators to be representable on the complement $C=[n]\setminus(A\cup B)$, yielding $k\leq|C|=n-2(d-1)$.
The resulting proof is short, elementary, and entirely algebraic. It avoids the use of von Neumann entropy, the no-cloning theorem, and the decoupling principle—tools that, while powerful, are not intrinsic to the stabiliser formalism.
Formal verification.
A secondary contribution of this work is a complete formalisation of the above argument in the Lean4 proof assistant, using the Mathlib library for finite-dimensional linear algebra. The formalisation covers the symplectic space model, the definition of stabiliser codes, the cleaning lemma, and the final bound, totalling approximately 1300 lines of Lean4 code. This is, to our knowledge, the first machine-checked proof of the Quantum Singleton Bound.
Formal verification of results in quantum information theory remains relatively rare. Most existing work in this area has focused on the verification of quantum programs and circuits rather than on coding-theoretic bounds. Tools such as SQIR [HRH+21], QHLProver [LZZ+19], CoqQ [ZBS+23], and the recent Veri-QEC framework [HFP+25] target operational verification—checking that a given circuit implements the intended unitary or that a specific error-correction procedure is correct. By contrast, our formalisation concerns a structural impossibility result about the parameters of any stabiliser code, which is a qualitatively different kind of claim. The linear-algebraic character of the proof makes it a natural candidate for mechanisation: the key steps are subspace dimension calculations and inclusions, which interact well with the finite-dimensional linear algebra API available in Mathlib. An entropy-based proof would require formalising the von Neumann entropy and its subadditivity properties, a substantially heavier undertaking.
2 Related work
Proofs of the Quantum Singleton Bound.
The Quantum Singleton Bound was first proved by Knill and Laflamme [KL97], who established the general theory of quantum error-correcting codes and the conditions under which errors can be corrected. The bound also appears in work of Rains [RAI99], who proved it (along with other bounds) using quantum weight enumerators and a linear programming approach. The textbook treatment in Nielsen and Chuang [NC10] derives the bound from the no-cloning theorem and entropy considerations. More recently, Grassl, Huber, and Winter [GHW22] provided a clean entropic proof and extended it to entanglement-assisted quantum error-correcting codes (EAQECCs) and catalytic codes (CQECCs), correcting a previously conjectured generalisation that turned out to be false [GRA21]. Klappenecker and Sarvepalli [KS07] proved that no $\mathbb{F}_{q}$-linear subsystem code over a prime field can beat the Quantum Singleton Bound.
The stabiliser formalism and symplectic geometry.
The symplectic vector space model of the Pauli group was introduced by Calderbank, Rains, Shor, and Sloane [CRS+97, CRS+98] and independently by Gottesman [GOT97]. In this framework, an $n$-qudit stabiliser code over $\mathbb{F}_{p}$ is described by an isotropic subspace of $\mathbb{F}_{p}^{2n}$ with respect to the standard symplectic form, and code properties such as distance and error correction translate into statements about subspace intersections and dimensions. The framework extends naturally to nonbinary (qudit) codes [AK01, KKK+06] and has been used to study the structure of logical operators [WIL09] and the synthesis of Clifford circuits via symplectic transformations [RCP+18].
The cleaning lemma.
The cleaning lemma, which asserts that logical operators on a correctable region can be “cleaned off” to the complementary region, has appeared in various forms across the literature. It is implicit in the work of Bravyi, Poulin, and Terhal on topological codes [BPT10] and is stated explicitly in Preskill’s lecture notes [PRE22], where a dimension-counting proof is given. Kalachev and Sadov [KS22] gave a systematic linear-algebraic treatment, showing that the cleaning lemma is at its core a fact about inner product spaces that admits a lattice-theoretic abstraction. The version we use here is closest to Preskill’s formulation, cast in symplectic language.
Formal verification of quantum information.
Formal verification of quantum computing results has seen growing interest. Hietala et al. [HRH+21] developed SQIR, a small quantum intermediate representation formalised in Coq, and used it to verify properties of quantum circuits including Grover’s algorithm. Liu et al. [LZZ+19] formalised quantum Hoare logic in Isabelle/HOL. Zhou et al. [ZBS+23] built CoqQ, a foundational Coq framework for verifying quantum programs based on the quantum-while language. Most recently, Huang et al. [HFP+25] proposed Veri-QEC, which formalises a program logic for quantum error correction in Coq and provides automated verification of fault-tolerant protocols. These efforts all target the verification of quantum programs or circuits. Our contribution is complementary: we formalise a coding-theoretic bound, which is a statement about the parameter space of all possible stabiliser codes rather than about any specific code or circuit.
@@ 3 Contributions
    We give a self-contained symplectic linear-algebraic proof of the Quantum Singleton Bound for $[[n,k,d]]$ stabiliser codes. The proof derives $k+2(d-1)\leq n$ from distance-based erasure correctability and the cleaning lemma via a dimension argument, without appealing to entropy or the no-cloning theorem.
@@
    We provide, to our knowledge, the first Lean4 formalisation of the Quantum Singleton Bound. The formalisation covers the full linear-algebraic development, including the symplectic form, stabiliser codes, the cleaning dimension identity, and the main theorem.
@@ 4 Conventions and symplectic model
with elements written $v=(v_{X},v_{Z})$. In the stabiliser formalism, a vector $v=(v_{X},v_{Z})\in V$ represents (the phase-free part of) the $n$-qudit Pauli operator $X^{v_{X}}Z^{v_{Z}}=\prod_{i=1}^{n}X_{i}^{v_{X}(i)}Z_{i}^{v_{Z}(i)}$, where $X_{i}$ and $Z_{i}$ are the generalised Pauli operators on the $i$-th qudit. Two Pauli operators commute if and only if the symplectic pairing of their corresponding vectors vanishes.
@@
This is a bilinear, alternating (and hence antisymmetric) form on $V$.
@@
Fix an index $i\in[n]$. Pairing $u$ with $(0,e_{i})$ gives $u_{X}(i)=0$, and pairing $u$ with $(e_{i},0)$ gives $u_{Z}(i)=0$. Since $i$ was arbitrary, all coordinates of $u$ vanish. ∎
Nondegeneracy ensures that the symplectic form identifies $V$ with its dual, which is the basis for the dimension identities used throughout the proof.
@@ 5 Support, restriction, and distance
Intuitively, $\mathrm{supp}(v)$ is the set of qudit positions on which the corresponding Pauli operator acts nontrivially.
@@
This is the subspace of $V$ consisting of vectors whose support is contained in $C$.
@@
Restriction to the coordinates indexed by $C$ gives a vector space isomorphism $V_{C}\cong\mathbb{F}^{|C|}\times\mathbb{F}^{|C|}$. ∎
@@
For $E\subseteq[n]$, define $r_{E}:V\to V_{E}$ by zeroing all coordinates outside $E$: $(r_{E}(v))_{X}(i)=v_{X}(i)$ and $(r_{E}(v))_{Z}(i)=v_{Z}(i)$ for $i\in E$, and zero otherwise.
@@
A stabiliser code $Q$ is specified by an isotropic subspace $S\subseteq V$ (i.e., $S\subseteq S^{\perp}$) with $\dim(S)=n-k$. The (Pauli) distance is
@@
Elements of $S$ represent stabiliser operators, and elements of $S^{\perp}\setminus S$ represent nontrivial logical operators. The quotient $S^{\perp}/S$ is the logical operator space.
@@
Nondegeneracy of the symplectic form gives $\dim(S)+\dim(S^{\perp})=2n$, so $\dim(S^{\perp})=n+k$. Therefore $\dim(S^{\perp}/S)=(n+k)-(n-k)=2k$. ∎
@@ 6 Erasure correctability
In the stabiliser setting, this condition is equivalent to the standard Knill–Laflamme error-correction conditions for the erasure channel that traces out the qudits in $E$ [KL97, NC10]. Concretely, $E$ is correctable if and only if no nontrivial logical operator is supported entirely within $E$: every element of $S^{\perp}$ that is supported on $E$ must already lie in the stabiliser $S$.
@@
Suppose $v\in S^{\perp}\cap V_{E}$ with $v\notin S$. Then $v\in S^{\perp}\setminus S$ and $\mathrm{wt}(v)\leq|E|\leq d-1$, contradicting the definition of $d$. ∎
@@ 7 Cleaning lemma and the dimension identity
The cleaning lemma captures a remarkable complementarity property of stabiliser codes: logical operators that “live” on a subset $M$ of qudit positions and those that live on the complementary set $M^{c}$ together account for all $2k$ independent logical degrees of freedom.
@@
This counts the number of independent logical Pauli operators (modulo stabilisers) that can be supported entirely on $M$.
@@
The following argument is a linear-algebraic reformulation of the dimension-counting proof appearing in Preskill’s notes [PRE22].
Let $P\cong V$ denote the abelianised Pauli group viewed as a $2n$-dimensional $\mathbb{F}$-vector space with the same symplectic form. Fix $M\subseteq[n]$ and write $M^{c}$ for its complement. Let $P_{M}:=V_{M}$ and $P_{M^{c}}:=V_{M^{c}}$.
Define the stabiliser-restriction subspaces
@@
Indeed, if a nontrivial linear combination of elements of $S_{M}\oplus S_{0}$ restricts to zero on $M$, then it is supported on $M^{c}$, hence lies in $S\cap P_{M^{c}}=S_{M^{c}}$, which is impossible unless the combination was trivial (since $S_{M}\oplus S_{M^{c}}\oplus S_{0}$ is a direct sum).
Now compute $\dim(S^{\perp}\cap P_{M})$. Since $P_{M}$ is symplectically orthogonal to $S_{M^{c}}$ (their supports are disjoint), the constraints for a vector in $P_{M}$ to commute with all of $S$ are exactly orthogonality to $r_{M}(S_{M}\oplus S_{0})=S_{M}\oplus r_{M}(S_{0})$. Therefore,
@@
The identity $g(M)+g(M^{c})=2k$ does not imply that each logical Pauli operator is supportable on exactly one of $M$ or $M^{c}$. It is a statement about dimensions, not about individual operators.
Corollary 1 (Cleaning lemma — support on the complement).
@@
If $M$ is correctable, then $S^{\perp}\cap V_{M}\subseteq S$, hence $g(M)=0$. By the dimension identity, $g(M^{c})=2k=\dim(S^{\perp}/S)$. ∎
8 Two disjoint correctable sets
The following lemma is the key step connecting the cleaning lemma to the Singleton bound. It shows that if we can find two disjoint correctable sets, then the number of logical qubits is bounded by the size of the remaining positions.
@@ 8 Two disjoint correctable sets
Set $D:=A^{c}=B\cup C$. Since $A$ is correctable we have $g(A)=0$, hence by the cleaning dimension identity (Lemma 5),
@@
so the quotient space $L:=(S^{\perp}\cap V_{D})/(S\cap V_{D})$ has dimension $\dim(L)=2k$.
@@
We claim $\bar{r}_{C}$ is injective. Suppose $\bar{r}_{C}([v])=0$. Then $r_{C}(v)\in r_{C}(W)$, so there exists $w\in W$ with $r_{C}(v)=r_{C}(w)$. Hence $r_{C}(v-w)=0$, so vwv-w is supported entirely on $B$, i.e., $v-w\in V_{B}$. Also $v-w\in S^{\perp}$ (since $v\in S^{\perp}$ and $w\in S$). Thus $v-w\in S^{\perp}\cap V_{B}$, and because $B$ is correctable we get $v-w\in S$. Since $w\in S$ as well, it follows that $v\in S$, hence $[v]=0$ in $L$.
@@
9 The Quantum Singleton Bound
@@
Choose disjoint subsets $A,B\subseteq[n]$ with |A|=|B|=d1|A|=|B|=d-1 and let $C=[n]\setminus(A\cup B)$. By Lemma 4, erasures of AA and $B$ are both correctable. Lemma 6 then yields
@@
If instead n<2(d1)n<2(d-1), then no such disjoint subsets exist. In this regime, the inequality $k\leq n-2(d-1)$ has a non-positive right-hand side, and since $k\geq 0$ for any stabiliser code, the bound is automatically satisfied. Hence the argument above covers all cases.
10 Discussion
We have derived the Quantum Singleton Bound from three ingredients: (i) the symplectic vector space model of stabiliser codes, (ii) distance-based erasure correctability, and (iii) a dimension-counting proof of the cleaning lemma following Preskill [PRE22]. The final inequality follows from a purely symplectic dimension argument using the correctability of two disjoint erasures.
Comparison with entropic proofs.
The entropy-based proof of the Quantum Singleton Bound [NC10, GHW22] proceeds by noting that if a code corrects the erasure of a set $E$, then the encoded information is fully determined by the complementary qudits, which yields $S(A)\geq S(AB)$ for appropriate subsystem decompositions (where $S$ denotes von Neumann entropy). The entropy approach is more general: it applies to arbitrary quantum codes, not only stabiliser codes, and it naturally handles approximate error correction and noisy settings. However, for the stabiliser case, the additional generality is not needed, and the symplectic proof makes the algebraic mechanism transparent.
Scope and limitations.
The proof presented here applies to stabiliser codes over prime fields $\mathbb{F}_{p}$. The extension to stabiliser codes over non-prime fields $\mathbb{F}_{q}=\mathbb{F}_{p^{m}}$ (using Hermitian or trace-symplectic forms) is straightforward and follows the same outline, as the relevant dimension identities remain valid in that setting [KKK+06, AK01]. For general (possibly non-additive) quantum codes, the entropic argument remains the most natural route.
The formalisation.
The Lean4 formalisation required several auxiliary developments not present in the current Mathlib library, including the symplectic form as a bilinear form, properties of support subspaces and restriction maps, and the interaction between symplectic orthogonal complements and subspace intersections. The most involved part of the formalisation was the cleaning dimension identity (Lemma 5), which required careful bookkeeping of the direct-sum decomposition $S=S_{M}\oplus S_{M^{c}}\oplus S_{0}$. The final theorem (Theorem 1) itself was a relatively short consequence once the lemmas were in place. We hope that the auxiliary material developed here will be useful for future formalisation efforts in quantum coding theory.